A dielectric with a positive charge of 13 C has 18.75X10 ^18 electrons added to it. What is the net charge of the dielectric in coulombs ?

1
Expert's answer
2021-11-09T13:26:04-0500

charge of electron:

"e=-1.6\cdot10^{-19}" C


net charge of the dielectric:

"Q=13-1.6\cdot10^{-19}\cdot 18.75\cdot10^{18}=13-3=10" C